A chemist has 1000 g of a solution that is 40% acid. How many grams of water must be added to reduce the acidity to 25%
Answer by stanbon(75887)

A chemist has 1000 g of a solution that is 40% acid. How many grams of water must be added to reduce the acidity to 25%
-----------------
Equation:
acid + acid = acid
0.40*1000 + 0*x = 0.25(1000+x)
------
Multiply thru by 100 to get:
40*1000 + 0 = 25*1000 + 25x
---
25x = 15*1000
x = (2/5)1000
x = 400 grapm(amt. of water needed)
=====================================
Cheers,
Stan H.
============